micheal is three times older than his sister angela. in 1 year micheal will be twice as old as angela. how old are the two children today?
Answer by Cromlix(4381)

micheal is three times older than his sister angela
M = 3A
in 1 year micheal will be twice as old as angela.
M + 1 = 2(A + 1)
Replace M with 3A from top equation.
3A + 1 = 2(A + 1)
3A + 1 = 2A + 2
Collect like terms.
3A - 2A = 2 - 1
A = 1
Current ages:
Michael = 3 years old
Angela = 1 year old.
Hope this helps.
:-)